There are two bags each containing red balls and blue balls. Bag A contains 1 red and 4 blue balls. Bag B contains 3 red and 13 blue balls. Lucky Louie says you should always choose Bag B if you are trying to draw a red ball because it has more red balls than Bag A. How do you respond ?

Answer by checkley77(12844)

1/(1+4)=1/5=.20=20% chance of picking a red ball from bag A.
3/(3+13)=1/16=.1875=18.75% chance of picking a red ballbag from bag B.
Bag A is a better bet.